OpenStudy (anonymous):

Solve for 4sec^2 x=8

OpenStudy (lgbasallote):

divide both sides by 4 first what do you have

OpenStudy (anonymous):

Sec^2/4 =8/4

OpenStudy (anonymous):

?? can you solve it i need it ASAP!!

OpenStudy (lgbasallote):

i can...but you need to..not me..im just gonna tell you how...it should be \[\sec^2 x = 8/4\] simplify 8/4

OpenStudy (apoorvk):

divide by 4. take square root. (don't forget '+' and '-' both) find corresponding inverse.